Loli:

Y=log x If y=10, then what is x? A. 10 B. 1 C. 100 D. 10^2

Nnesha:

`y=log x` and `y=10` set both equation equal to each othr \[\huge\rm 10=\log x\] then rewrite this log equation to in an exponential form \[\large\rm \log_\color{Red}{b} \color{orange}{x} =\color{blue}{y} \] \[\large\rm \color{orange}{x}= \color{red}{b}^\color{blue}{y}\]

Nnesha:

\[\log = \log_{10}\]

Nnesha:

options aren't correct or idk how to do math anymore rip mai life

satellite73:

options are wrong or there is a typo \[10=\log(x)\iff x=10^{10}\]

AndreaTheShorty:

I believe the answer is 10^2.

Nnesha:

another way to find the answer is the plugin each option for x \[10 = \log (10^2)\] z left side equal to right side ? if not then that's the wrong answer.

Nnesha:

to*

HuskyNation:

It would be D ^_^

Nnesha:

prove it